# Probabilities of Chat LLMs Are Miscalibrated but Still Predict Correctness on Multiple-Choice Q&A

### v3

@@ Appendix F Dataset-level results
Falcon 7B 50.1 49.4 0/2 50.4 0/2
Falcon 40B 50.8 52.5 1/2 51.5 1/2
Llama 2 7B 51.0 50.8 0/2 50.9 0/2
Llama 2 70B 51.9 55.5 2/2 51.6 0/2
Llama 3.0 8B 55.8 56.8 2/2 56.5 2/2
Llama 3.0 70B 74.9 74.0 2/2 61.0 2/2
Llama 3.1 8B 57.3 57.6 2/2 55.7 2/2
Llama 3.1 70B 77.8 75.8 2/2 59.7 2/2
Mistral 7B 50.8 59.5 2/2 59.5 2/2
Mixtral 8x7B 62.8 53.4 1/2 54.4 2/2
SOLAR 10.7B 68.8 54.7 2/2 57.4 2/2
Yi 6B 58.7 56.1 2/2 55.2 2/2
Yi 34B 67.5 49.2 1/2 50.7 2/2
GPT-3.5 Turbo 60.8 61.8 2/2 2/2
GPT-4o 81.9 78.3 2/2 2/2
Table 12: AUROC results for WinoGrande. See Table 1 for more explanation.
@@
LLM Base MSP Max Logit Base MSP Max Logit
Falcon 7B $-48.6$ $-0.9$ $-0.3$ $-122.9$ $-2.2$ $-0.9$
Falcon 40B $5.9$ $13.7$ $13.4$ $-41.1$ $-9.3$ $-6.7$
Llama 2 7B $-9.6$ $-5.2$ $0.2$ $-64.4$ $-10.5$ $-26.2$
Llama 2 70B $48.4$ $48.4$ $48.4$ $22.6$ $39.6$ $24.6$
Llama 3.0 8B $49.3$ $53.4$ $50.0$ $23.9$ $40.2$ $34.2$
Llama 3.0 70B $84.7$ $84.7$ $84.7$ $77.0$ $77.0$ $77.0$
Llama 3.1 8B $54.7$ $57.5$ $54.6$ $32.0$ $45.3$ $36.8$
Llama 3.1 70B $86.8$ $86.8$ $86.8$ $80.2$ $80.2$ $80.2$
Mistral 7B $41.3$ $35.4$ $36.2$ $12.0$ $21.8$ $21.8$
Mixtral 8x7B $68.3$ $68.3$ $68.3$ $52.5$ $52.5$ $52.5$
SOLAR 10.7B $60.9$ $60.9$ $62.0$ $41.4$ $41.3$ $46.7$
Yi 6B $37.6$ $41.1$ $31.2$ $6.3$ $26.1$ $10.9$
Yi 34B $70.7$ $71.2$ $70.7$ $56.1$ $58.3$ $55.5$
GPT-3.5 Turbo $66.7$ $67.6$ $50.0$ $58.7$
GPT-4o $92.6$ $92.6$ $88.9$ $88.9$
Table 13: Q&A with abstention results for ARC-Challenge. See Table 2 for an explanation of the scoring scheme.
@@
LLM Base MSP Max Logit Base MSP Max Logit
Falcon 7B 00 $98.2$ $99.2$ 00 $98.2$ $99.2$
Falcon 40B 00 $40.1$ $26.8$ 00 $40.1$ $85.2$
Llama 2 7B 00 $10.7$ $47.1$ 00 $63.7$ $47.1$
Llama 2 70B 00 $0.0$ $0.0$ 00 $28.9$ $57.0$
Llama 3.0 8B 00 $20.3$ $18.4$ 00 $20.3$ $18.4$
Llama 3.0 70B 00 $0.0$ $0.0$ 00 $0.0$ $0.0$
Llama 3.1 8B 00 $18.1$ $9.8$ 00 $18.1$ $9.8$
Llama 3.1 70B 00 $0.0$ $0.0$ 00 $0.0$ $0.0$
Mistral 7B 00 $37.2$ $35.0$ 00 $37.2$ $35.0$
Mixtral 8x7B 00 $0.0$ $0.0$ 00 $0.0$ $0.0$
SOLAR 10.7B 00 $0.0$ $7.5$ 00 $7.1$ $7.5$
Yi 6B 00 $18.7$ $28.1$ 00 $30.8$ $28.1$
Yi 34B 00 $2.9$ $0.0$ 00 $2.9$ $8.1$
GPT-3.5 Turbo 00 $6.0$ 00 $23.7$
GPT-4o 00 $0.0$ 00 $0.0$
Table 14: Frequency of abstention on ARC-Challenge in the Section 6 experiments.
@@
LLM Base MSP Max Logit Base MSP Max Logit
Falcon 7B $-50.2$ $-11.1$ $-0.7$ $-125.3$ $-4.2$ $-1.8$
Falcon 40B $-14.3$ $0.8$ $0.5$ $-71.4$ $-0.3$ $0.2$
Llama 2 7B $-17.0$ $-0.4$ $-1.1$ $-75.5$ $-12.6$ $-5.0$
Llama 2 70B $27.8$ $32.0$ $30.3$ $-8.3$ $7.7$ $10.6$
Llama 3.0 8B $31.2$ $35.4$ $32.0$ $-3.1$ $18.3$ $17.2$
Llama 3.0 70B $55.3$ $55.3$ $54.5$ $32.9$ $38.2$ $38.4$
Llama 3.1 8B $32.2$ $33.0$ $32.2$ $-1.7$ $20.5$ $14.4$
Llama 3.1 70B $57.9$ $58.4$ $57.9$ $36.8$ $46.9$ $36.8$
Mistral 7B $5.0$ $11.1$ $10.2$ $-42.5$ $-7.5$ $2.0$
Mixtral 8x7B $31.6$ $31.9$ $826$-2.6-2.6- 2.6 $4.0$ $7.1$
SOLAR 10.7B $58.4$ $58.4$ $58.4$ $37.6$ $41.5$ $37.6$
Yi 6B $-15.9$ $4.5$ $0.7$ $-73.9$ $1.8$ $-0.1$
Yi 34B $50.9$ $51.1$ $49.9$ $26.3$ $28.5$ $32.8$
GPT-3.5 Turbo $45.4$ $47.2$ $18.1$ $32.9$
GPT-4o $76.9$ $76.9$ $65.3$ $65.3$
Table 15: Q&A with abstention results for HellaSwag. See Table 2 for an explanation of the scoring scheme.
@@
LLM Base MSP Max Logit Base MSP Max Logit
Falcon 7B 00 $78.2$ $98.5$ 00 $96.7$ $98.5$
Falcon 40B 00 $97.2$ $98.9$ 00 $97.2$ $98.9$
Llama 2 7B 00 $75.9$ $93.2$ 00 $75.9$ $93.2$
Llama 2 70B 00 $19.4$ $13.9$ 00 $19.4$ $56.9$
Llama 3.0 8B 00 $30.6$ $7.1$ 00 $30.6$ $55.0$
Llama 3.0 70B 00 $0.0$ $13.2$ 00 $6.8$ $13.2$
Llama 3.1 8B 00 $1.5$ $0.0$ 00 $61.6$ $55.6$
Llama 3.1 70B 00 $1.5$ $0.0$ 00 $39.3$ $0.0$
Mistral 7B 00 $51.6$ $64.1$ 00 $51.6$ $86.5$
Mixtral 8x7B 00 $12.3$ $33.8$ 00 $12.3$ $33.8$
SOLAR 10.7B 00 $0.0$ $0.0$ 00 $8.4$ $0.0$
Yi 6B 00 $90.1$ $97.8$ 00 $90.1$ $97.8$
Yi 34B 00 $0.7$ $11.2$ 00 $56.3$ $28.0$
GPT-3.5 Turbo 00 $9.6$ 00 $26.3$
GPT-4o 00 $0.0$ 00 $0.0$
Table 16: Frequency of abstention on HellaSwag in the Section 6 experiments.
Balanced Conservative
LLM Base MSP Max Logit Base MSP Max Logit
Falcon 7B $-48.3$ $-0.5$ $0.0$ $-122.4$ $-1.3$ $0.0$
Falcon 40B $-12.3$ $-6.6$ $-7.6$ $-68.4$ $-0.6$ $-20.6$
Llama 2 7B $-19.6$ $3.0$ $-13.5$ $-79.5$ $-1.4$ $-0.2$
Llama 2 70B $12.1$ $18.5$ $13.0$ $-31.9$ $-13.2$ $5.1$
Llama 3.0 8B $12.5$ $12.8$ $12.5$ $-31.3$ $16.4$ $9.8$
Llama 3.0 70B $49.9$ $53.2$ $49.9$ $24.9$ $43.7$ $41.3$
Llama 3.1 8B $16.3$ $27.6$ $19.4$ $-25.6$ $21.2$ $16.8$
Llama 3.1 70B $59.6$ $61.4$ $59.6$ $39.5$ $46.7$ $42.9$
Mistral 7B $5.0$ $10.8$ $7.9$ $-42.6$ $-24.5$ $-34.5$
Mixtral 8x7B $30.4$ $31.6$ $32.9$ $-4.5$ $0.8$ $10.0$
SOLAR 10.7B $16.0$ $16.0$ $17.5$ $-26.0$ $-26.0$ $-12.2$
Yi 6B $3.8$ $15.4$ $7.8$ $-44.3$ $6.7$ $-17.5$
Yi 34B $31.1$ $31.1$ $31.1$ $-3.4$ $10.6$ $-3.4$
GPT-3.5 Turbo $30.8$ $36.0$ $-3.7$ $13.7$
GPT-4o $67.9$ $68.0$ $51.9$ $61.9$
Table 17: Q&A with abstention results for MMLU. See Table 2 for an explanation of the scoring scheme.
Balanced Conservative
LLM Base MSP Max Logit Base MSP Max Logit
Falcon 7B 00 $98.9$ $100.0$ 00 $98.9$ $100.0$
Falcon 40B 00 $11.3$ $7.7$ 00 $93.3$ $69.8$
Llama 2 7B 00 $86.2$ $13.6$ 00 $95.4$ $99.0$
Llama 2 70B 00 $18.2$ $4.0$ 00 $18.2$ $68.0$
Llama 3.0 8B 00 $0.9$ $0.1$ 00 $55.9$ $51.8$
Llama 3.0 70B 00 $9.6$ $0.0$ 00 $27.2$ $34.9$
Llama 3.1 8B 00 $59.8$ $8.0$ 00 $59.8$ $74.7$
Llama 3.1 70B 00 $9.2$ $0.0$ 00 $9.2$ $26.6$
Mistral 7B 00 $18.6$ $7.4$ 00 $18.6$ $7.4$
Mixtral 8x7B 00 $6.9$ $14.9$ 00 $6.9$ $22.5$
SOLAR 10.7B 00 $0.0$ $4.9$ 00 $0.0$ $15.4$
Yi 6B 00 $41.8$ $41.5$ 00 $65.4$ $41.5$
Yi 34B 00 $0.0$ $0.0$ 00 $35.3$ $0.0$
GPT-3.5 Turbo 00 $19.3$ 00 $19.3$
GPT-4o 00 $0.7$ 00 $17.1$
Table 18: Frequency of abstention on MMLU in the Section 6 experiments.
Balanced Conservative
LLM Base MSP Max Logit Base MSP Max Logit
Falcon 7B $-59.1$ $-11.3$ $-2.2$ $-138.7$ $-10.3$ $-5.0$
Falcon 40B $-42.6$ $-3.0$ $-0.7$ $-113.9$ $-7.9$ $-2.5$
Llama 2 7B $-52.4$ $-39.7$ $-44.4$ $-128.6$ $-4.7$ $-0.6$
Llama 2 70B $-11.2$ $8.7$ $4.0$ $-66.7$ $-16.8$ $-19.6$
Llama 3.0 8B $-19.5$ $-5.0$ $1.4$ $-79.2$ $-0.9$ $-9.1$
Llama 3.0 70B $44.2$ $46.4$ $39.7$ $16.4$ $38.9$ $28.3$
Llama 3.1 8B $10.4$ $22.2$ $11.7$ $-34.4$ $9.8$ $-4.6$
Llama 3.1 70B $48.2$ $49.3$ $48.2$ $22.3$ $41.8$ $26.3$
Mistral 7B $10.3$ $17.6$ $12.5$ $-34.6$ $-8.9$ $3.3$
Mixtral 8x7B $34.7$ $37.3$ $32.9$ $2.0$ $6.8$ $14.7$
SOLAR 10.7B $-2.0$ $4.9$ $8.9$ $-53.0$ $-8.6$ $-8.1$
Yi 6B $-12.8$ $-7.9$ $-5.9$ $-69.1$ $-4.6$ $-12.1$
Yi 34B $13.0$ $14.5$ $14.5$ $-30.5$ $-0.2$ $-25.6$
GPT-3.5 Turbo $10.4$ $21.6$ $-34.4$ $14.3$
GPT-4o $64.9$ $65.7$ $47.4$ $55.9$
Table 19: Q&A with abstention results for TruthfulQA. See Table 2 for an explanation of the scoring scheme.
Balanced Conservative
LLM Base MSP Max Logit Base MSP Max Logit
Falcon 7B 00 $72.6$ $96.8$ 00 $89.1$ $96.8$
Falcon 40B 00 $93.4$ $97.1$ 00 $93.4$ $97.1$
Llama 2 7B 00 $21.0$ $12.3$ 00 $95.8$ $99.3$
Llama 2 70B 00 $40.4$ $34.8$ 00 $40.4$ $45.6$
Llama 3.0 8B 00 $28.9$ $66.0$ 00 $81.6$ $73.4$
Llama 3.0 70B 00 $38.6$ $37.7$ 00 $38.6$ $37.7$
Llama 3.1 8B 00 $33.1$ $52.2$ 00 $68.6$ $56.8$
Llama 3.1 70B 00 $38.8$ $0.0$ 00 $46.5$ $9.4$
Mistral 7B 00 $24.8$ $11.4$ 00 $30.0$ $81.1$
Mixtral 8x7B 00 $9.2$ $29.2$ 00 $79.2$ $39.8$
SOLAR 10.7B 00 $43.3$ $51.7$ 00 $71.8$ $61.4$
Yi 6B 00 $12.5$ $19.8$ 00 $72.7$ $71.6$
Yi 34B 00 $3.5$ $5.3$ 00 $47.5$ $5.3$
GPT-3.5 Turbo 00 $63.6$ 00 $63.6$
GPT-4o 00 $16.0$ 00 $18.0$
Table 20: Frequency of abstention on TruthfulQA in the Section 6 experiments.
Balanced Conservative
LLM Base MSP Max Logit Base MSP Max Logit
Falcon 7B $0.1$ $0.1$ $0.1$ $-49.8$ $-8.7$ $-4.1$
Falcon 40B $1.6$ $1.1$ $2.3$ $-47.6$ $-6.2$ $-6.3$
Llama 2 7B $1.9$ $1.9$ $1.9$ $-47.1$ $-2.5$ $-0.5$
Llama 2 70B $3.8$ $6.8$ $2.9$ $-44.3$ $-15.9$ $-20.8$
Llama 3.0 8B $11.7$ $11.4$ $6.4$ $-32.5$ $-3.5$ $1.0$
Llama 3.0 70B $49.7$ $46.9$ $47.0$ $24.6$ $34.0$ $25.9$
Llama 3.1 8B $14.7$ $14.6$ $14.7$ $-27.9$ $-0.6$ $0.7$
Llama 3.1 70B $55.7$ $54.2$ $51.9$ $33.5$ $40.4$ $27.3$
Mistral 7B $1.7$ $5.8$ $6.0$ $-47.5$ $0.0$ $-1.0$
Mixtral 8x7B $25.6$ $25.6$ $24.0$ $-11.6$ $-0.6$ $-1.8$
SOLAR 10.7B $37.6$ $34.5$ $37.6$ $6.4$ $6.9$ $6.4$
Yi 6B $17.4$ $17.4$ $17.4$ $-23.9$ $0.2$ $0.0$
Yi 34B $35.0$ $35.0$ $35.0$ $2.5$ $2.5$ $2.5$
GPT-3.5 Turbo $21.6$ $16.9$ $-17.7$ $5.9$
GPT-4o $63.8$ $63.8$ $45.8$ $45.8$
Table 21: Q&A with abstention results for WinoGrande. See Table 2 for an explanation of the scoring scheme.
Balanced Conservative
LLM Base MSP Max Logit Base MSP Max Logit
Falcon 7B 00 $0.0$ $0.0$ 00 $82.5$ $91.0$
Falcon 40B 00 $84.2$ $20.7$ 00 $84.2$ $85.0$
Llama 2 7B 00 $0.0$ $0.0$ 00 $93.7$ $98.5$
Llama 2 70B 00 $15.2$ $49.8$ 00 $50.1$ $49.8$
Llama 3.0 8B 00 $21.5$ $76.0$ 00 $67.7$ $93.5$
Llama 3.0 70B 00 $27.2$ $10.7$ 00 $27.2$ $10.7$
Llama 3.1 8B 00 $1.5$ $0.0$ 00 $68.8$ $99.0$
Llama 3.1 70B 00 $16.4$ $9.8$ 00 $22.1$ $33.4$
Mistral 7B 00 $39.8$ $41.9$ 00 $100.0$ $80.9$
Mixtral 8x7B 00 $0.0$ $13.4$ 00 $78.3$ $49.6$
SOLAR 10.7B 00 $10.1$ $0.0$ 00 $10.1$ $0.0$
Yi 6B 00 $0.0$ $0.0$ 00 $98.8$ $100.0$
Yi 34B 00 $0.0$ $0.0$ 00 $0.0$ $0.0$
GPT-3.5 Turbo 00 $61.1$ 00 $61.1$
GPT-4o 00 $0.0$ 00 $0.0$
Table 22: Frequency of abstention on WinoGrande in the Section 6 experiments.

### v4

@@ E.2 HellaSwag
Falcon 7B 25.0 50.3 0/2 49.9 0/2
Falcon 40B 30.4 55.0 1/2 60.5 2/2
Llama 2 7B 38.7 60.3 2/2 57.8 2/2
Llama 2 70B 67.3 71.8 2/2 69.7 2/2
Llama 3.0 8B 65.1 72.4 2/2 69.8 2/2
Llama 3.0 70B 78.7 82.3 2/2 68.3 2/2
Llama 3.1 8B 64.4 72.4 2/2 67.2 2/2
Llama 3.1 70B 78.4 82.1 2/2 61.9 2/2
Mistral 7B 52.5 62.7 2/2 62.4 2/2
Mixtral 8x7B 69.0 68.3 2/2 58.3 2/2
SOLAR 10.7B 80.9 76.4 2/2 72.1 2/2
Yi 6B 35.0 68.0 2/2 62.7 2/2
Yi 34B 72.1 75.5 2/2 70.3 2/2
GPT-3.5 Turbo 71.0 77.8 2/2
GPT-4o 88.3 89.7 2/2
Figure 23: AUROC results for HellaSwag for MSP (left) and Max Logit (right). Both MSP AUROC and Max Logit AUROC exhibit significant correlations with Q&A accuracy: $R^{2}=0.85$ ($p<10^{-4}$) and $R^{2}=0.52$ ($p=0.005$), respectively.
Table 20: Q&A-with-abstention results for HellaSwag. See Table 4 for an explanation of the scoring scheme.
@@
LLM No abstain MSP Max Logit No abstain MSP Max Logit
Falcon 7B 49.9-49.9 1.3-1.3 14.4-14.4 124.9-124.9 3.6-3.6 36.2-36.2
Falcon 40B 39.2-39.2 1.0-1.0 $0.1$ 108.9-108.9 5.6-5.6 4.9-4.9
Llama 2 7B 22.6-22.6 $1.5$ $0.1$ 83.9-83.9 1.8-1.8 $0.0$
Llama 2 70B $34.7$ $35.9$ $34.9$ $2.0$ $21.8$ $16.7$
Llama 3.0 8B $30.3$ $34.4$ $30.3$ 4.6-4.6 $18.1$ $13.6$
Llama 3.0 70B $57.4$ $57.4$ $56.2$ $36.1$ $43.2$ $38.6$
Llama 3.1 8B $28.8$ $32.1$ $29.5$ 6.8-6.8 $10.5$ $10.1$
Llama 3.1 70B $56.8$ $57.6$ $56.8$ $35.2$ $38.8$ $33.1$
Mistral 7B $5.1$ $6.6$ $10.0$ 42.4-42.4 $0.0$ 4.9-4.9
Mixtral 8x7B $38.1$ $39.7$ $12.9$ $7.1$ $5.6$ $6.0$
SOLAR 10.7B $61.7$ $61.7$ $61.7$ $42.6$ $42.6$ $42.6$
Yi 6B 30.0-30.0 $2.9$ $0.4$ 95.1-95.1 $1.2$ $0.0$
Yi 34B $44.3$ $45.8$ $37.0$ $16.4$ $31.1$ $23.8$
GPT-3.5 Turbo $42.0$ $43.6$ $13.1$ $32.3$
GPT-4o $76.6$ $76.6$ $64.9$ $64.9$
Figure 24: Q&A-with-abstention scores for HellaSwag as a function of the amount of training data. The x-axis is the number of data points included in the training data (referred to as $k$ in Section˜6.1).
E.3 MMLU
Figure 25: Calibration results for MMLU. The coefficient of determination between Q&A accuracy and calibration error was $R^{2}=0.03$ ($p=0.57)$.
Table 21: AUROC results for MMLU. See Table 2 for an explanation of the $p$-values.
MSP Max Logit
LLM Q&A accuracy AUROC $p<10^{-4}$ AUROC $p<10^{-4}$
Falcon 7B 25.6 51.1 0/2 50.7 0/2
Falcon 40B 41.6 66.0 2/2 60.3 2/2
Llama 2 7B 39.4 64.5 2/2 61.7 2/2
Llama 2 70B 55.8 72.7 2/2 69.8 2/2
Llama 3.0 8B 56.3 77.0 2/2 76.1 2/2
Llama 3.0 70B 74.8 83.0 2/2 74.6 2/2
Llama 3.1 8B 57.9 77.7 2/2 73.4 2/2
Llama 3.1 70B 79.3 84.3 2/2 75.1 2/2
Mistral 7B 53.1 68.9 2/2 64.8 2/2
Mixtral 8x7B 65.7 74.8 2/2 65.3 2/2
SOLAR 10.7B 57.4 70.9 2/2 65.7 2/2
Yi 6B 48.9 68.3 2/2 61.8 2/2
Yi 34B 62.8 75.2 2/2 69.7 2/2
GPT-3.5 Turbo 65.5 80.1 2/2
GPT-4o 84.2 85.6 2/2
Figure 26: AUROC results for MMLU for MSP (left) and Max Logit (right). Both MSP AUROC and Max Logit AUROC exhibit significant correlations with Q&A accuracy: $R^{2}=0.92$ ($p<10^{-4}$) and $R^{2}=0.72$ ($p<10^{-3}$), respectively.
Table 22: Q&A-with-abstention results for MMLU. See Table 4 for an explanation of the scoring scheme.
@@ E.3 MMLU
LLM No abstain MSP Max Logit No abstain MSP Max Logit
Falcon 7B 48.7-48.7 3.2-3.2 2.7-2.7 123.1-123.1 1.0-1.0 7.0-7.0
Falcon 40B 16.9-16.9 1.2-1.2 11.0-11.0 75.3-75.3 $3.0$ $0.2$
Llama 2 7B 21.3-21.3 $6.2$ $2.8$ 81.9-81.9 $1.3$ 3.5-3.5
Llama 2 70B $11.5$ $11.5$ $11.5$ 32.7-32.7 $13.8$ 32.7-32.7
Llama 3.0 8B $12.6$ $14.2$ $14.7$ 31.0-31.0 $18.1$ $14.2$
Llama 3.0 70B $49.6$ $49.6$ $49.6$ $24.4$ $43.7$ $24.4$
Llama 3.1 8B $15.9$ $27.9$ $18.5$ 26.2-26.2 $19.9$ 17.9-17.9
Llama 3.1 70B $58.6$ $61.2$ $58.6$ $37.9$ $47.5$ $43.1$
Mistral 7B $6.1$ $6.1$ $9.4$ 40.9-40.9 20.8-20.8 30.1-30.1
Mixtral 8x7B $31.2$ $31.2$ $31.4$ 3.1-3.1 3.1-3.1 2.6-2.6
SOLAR 10.7B $14.9$ $21.5$ $20.6$ 27.6-27.6 $0.0$ $5.5$
Yi 6B 2.2-2.2 $11.2$ $0.9$ 53.4-53.4 $3.6$ 3.7-3.7
Yi 34B $25.5$ $31.1$ $29.3$ 11.8-11.8 $7.9$ $14.5$
GPT-3.5 Turbo $31.0$ $38.1$ 3.5-3.5 $30.1$
GPT-4o $68.3$ $68.5$ $52.5$ $53.0$
Figure 27: Q&A-with-abstention scores for MMLU as a function of the amount of training data. The x-axis is the number of data points included in the training data (referred to as $k$ in Section˜6.1).
E.4 TruthfulQA
Figure 28: Calibration results for TruthfulQA. The coefficient of determination between Q&A accuracy and calibration error was $R^{2}=0.17$ ($p=0.13)$.
Table 23: AUROC results for TruthfulQA. See Table 2 for an explanation of the $p$-values.
MSP Max Logit
LLM Q&A accuracy AUROC $p<10^{-4}$ AUROC $p<10^{-4}$
Falcon 7B 22.7 56.5 1/2 57.2 1/2
Falcon 40B 31.9 59.3 2/2 55.0 1/2
Llama 2 7B 25.4 53.3 0/2 57.2 1/2
Llama 2 70B 46.3 72.2 2/2 67.5 2/2
Llama 3.0 8B 42.5 69.1 2/2 62.9 2/2
Llama 3.0 70B 72.4 78.6 2/2 66.8 2/2
Llama 3.1 8B 53.9 74.0 2/2 62.2 2/2
Llama 3.1 70B 73.9 84.1 2/2 67.7 2/2
Mistral 7B 55.2 71.1 2/2 63.4 2/2
Mixtral 8x7B 68.7 73.8 2/2 62.7 2/2
SOLAR 10.7B 52.4 72.0 2/2 70.0 2/2
Yi 6B 40.0 65.8 2/2 61.2 2/2
Yi 34B 61.0 77.5 2/2 68.5 2/2
GPT-3.5 Turbo 55.9 75.3 2/2
GPT-4o 84.0 82.9 2/2
Figure 29: AUROC results for TruthfulQA for MSP (left) and Max Logit (right). Both MSP AUROC and Max Logit AUROC exhibit significant correlations with Q&A accuracy: $R^{2}=0.90$ ($p<10^{-4}$) and $R^{2}=0.56$ ($p=0.004$), respectively.
Table 24: Q&A-with-abstention results for TruthfulQA. See Table 4 for an explanation of the scoring scheme.
@@ E.4 TruthfulQA
LLM No abstain MSP Max Logit No abstain MSP Max Logit
Falcon 7B 54.8-54.8 17.0-17.0 20.9-20.9 132.2-132.2 23.7-23.7 14.4-14.4
Falcon 40B 36.2-36.2 0.7-0.7 4.5-4.5 104.3-104.3 1.0-1.0 11.9-11.9
Llama 2 7B 49.6-49.6 29.1-29.1 3.0-3.0 124.3-124.3 6.3-6.3 6.1-6.1
Llama 2 70B 7.4-7.4 $14.1$ $4.0$ 61.2-61.2 0.5-0.5 0.6-0.6
Llama 3.0 8B 15.2-15.2 2.5-2.5 $1.7$ 72.9-72.9 $1.4$ 8.2-8.2
Llama 3.0 70B $44.9$ $46.1$ $44.9$ $17.3$ $37.4$ $18.8$
Llama 3.1 8B $7.7$ $21.9$ $9.6$ 38.5-38.5 $5.8$ 2.2-2.2
Llama 3.1 70B $48.0$ $49.4$ $48.0$ $21.9$ $38.9$ $26.6$
Mistral 7B $10.2$ $18.9$ $11.0$ 34.8-34.8 9.9-9.9 $2.0$
Mixtral 8x7B $37.3$ $40.0$ $37.3$ $5.9$ $13.1$ $14.9$
SOLAR 10.7B $4.7$ $8.4$ $17.7$ 42.9-42.9 33.9-33.9 $7.1$
Yi 6B 20.4-20.4 6.6-6.6 17.8-17.8 80.7-80.7 $1.4$ 26.7-26.7
Yi 34B $21.7$ $22.1$ $26.7$ 17.5-17.5 $24.0$ $4.8$
GPT-3.5 Turbo $11.8$ $24.5$ 32.3-32.3 $11.6$
GPT-4o $68.2$ $65.3$ $52.2$ $58.1$
Figure 30: Q&A-with-abstention scores for TruthfulQA as a function of the amount of training data. The x-axis is the number of data points included in the training data (referred to as $k$ in Section˜6.1).
E.5 WinoGrande
Figure 31: Calibration results for WinoGrande. The coefficient of determination between Q&A accuracy and calibration error was $R^{2}=0.10$ ($p=0.25)$.
Table 25: AUROC results for WinoGrande. See Table 2 for an explanation of the $p$-values.
MSP Max Logit
LLM Q&A accuracy AUROC $p<10^{-4}$ AUROC $p<10^{-4}$
Falcon 7B 50.2 49.9 0/2 50.6 0/2
Falcon 40B 50.9 52.0 0/2 50.1 0/2
Llama 2 7B 51.4 51.1 0/2 50.3 0/2
Llama 2 70B 50.4 60.1 2/2 53.2 1/2
Llama 3.0 8B 55.6 55.8 2/2 56.4 2/2
Llama 3.0 70B 75.3 73.3 2/2 54.6 1/2
Llama 3.1 8B 56.9 58.1 2/2 55.8 2/2
Llama 3.1 70B 78.2 75.9 2/2 63.9 2/2
Mistral 7B 53.6 55.9 2/2 54.3 2/2
Mixtral 8x7B 62.9 54.2 2/2 50.1 0/2
SOLAR 10.7B 67.1 63.0 2/2 55.3 2/2
Yi 6B 58.3 57.8 2/2 55.7 2/2
Yi 34B 67.2 66.0 2/2 62.4 2/2
GPT-3.5 Turbo 61.7 62.8 2/2
GPT-4o 81.7 76.4 2/2
Figure 32: AUROC results for WinoGrande for MSP (left) and Max Logit (right). MSP AUROC exhibits a significant correlation with Q&A accuracy ($R^{2}=0.86,p<10^{-4})$. However, the correlation between Max Logit AUROC and Q&A accuracy is much weaker ($R^{2}=0.45,p<10^{-3}$).
Table 26: Q&A-with-abstention results for WinoGrande. See Table 4 for an explanation of the scoring scheme.
@@ E.5 WinoGrande
LLM No abstain MSP Max Logit No abstain MSP Max Logit
Falcon 7B $0.2$ $0.1$ $0.2$ 49.7-49.7 39.6-39.6 29.7-29.7
Falcon 40B $1.8$ $1.8$ $1.1$ 47.3-47.3 1.7-1.7 19.8-19.8
Llama 2 7B $2.7$ $2.7$ $2.7$ 45.9-45.9 8.8-8.8 12.7-12.7
Llama 2 70B $0.8$ $7.4$ $0.8$ 48.9-48.9 5.5-5.5 0.9-0.9
Llama 3.0 8B $11.3$ $11.1$ $6.3$ 33.1-33.1 $1.1$ 0.1-0.1
Llama 3.0 70B $50.7$ $49.7$ $47.5$ $26.0$ $31.0$ $21.7$
Llama 3.1 8B $13.8$ $14.0$ $8.6$ 29.2-29.2 $1.8$ 6.5-6.5
Llama 3.1 70B $56.3$ $56.2$ $54.1$ $34.5$ $42.0$ $34.6$
Mistral 7B $7.2$ $0.0$ $6.3$ 39.2-39.2 $0.0$ $0.7$
Mixtral 8x7B $25.8$ $25.8$ $25.0$ 11.3-11.3 11.3-11.3 10.2-10.2
SOLAR 10.7B $34.2$ $34.2$ $34.2$ $1.3$ $1.3$ $1.3$
Yi 6B $16.5$ $15.4$ $10.6$ 25.2-25.2 $1.6$ $0.5$
Yi 34B $34.3$ $33.4$ $32.2$ $1.4$ $9.4$ $6.6$
GPT-3.5 Turbo $23.4$ $23.4$ 15.0-15.0 $7.7$
GPT-4o $63.3$ $63.3$ $45.0$ $45.0$
Figure 33: Q&A-with-abstention scores for WinoGrande as a function of the amount of training data. The x-axis is the number of data points included in the training data (referred to as $k$ in Section˜6.1).